# Exactly solvable model of the square-root price impact dynamics under the long-range market-order correlation

### v2

@@ Abstract
In econophysics, there are several enigmatic empirical laws: (i) the market-order flow has strong persistence (long-range order-sign correlation), well formulated as the Lillo-Mike-Farmer model. This phenomenon seems paradoxical given the diffusive and unpredictable price dynamics; (ii) the price impact $I(Q)$ of a large metaorder $Q$ follows the square-root law, $I(Q)\propto\sqrt{Q}$. In this Letter, we propose an exactly solvable model of the nonlinear price-impact dynamics that unifies these enigmas. We generalize the Lillo-Mike-Farmer model to nonlinear price-impact dynamics, which is mapped to an exactly solvable Lévy-walk model. Our exact solution and numerical simulations reveal three important points: First, the price dynamics remains diffusive under the square-root law, even under the long-range correlation. Second, price-movement statistics follows truncated power laws with typical exponent around three. Third, volatility has long memory. While this simple model lacks adjustable free parameters, it naturally aligns even with other enigmatic empirical laws, such as (iii) the inverse-cubic law for price statistics and (iv) volatility clustering. This work illustrates the crucial role of the square-root law in understanding rich and complex financial price dynamics from a single coherent viewpoint.
econophysics, power-law scalings, the square-root law, the Lillo-Mike-Farmer model, Lévy walks
@@ Introduction.
In econophysics, there are four empirical enigmas [1, 2, 3]. The first enigma is the long-range correlation (LRC) of the market-order flow. Market order is an immediate decision to buy or sell stocks at the best prices. By representing the buy (sell) market order at time $t$ by $\epsilon_{t}=+1$ ($\epsilon_{t}=-1$), the LRC states that the autocorrelation (ACF) of the market-order signs exhibits a long memory
$\langle\epsilon_{t}\epsilon_{t+\tau}\rangle_{\rm ss}\propto\tau^{-\gamma},\>\>\>\gamma\in(0,1),$ (1)
where $\langle\dots\rangle_{\rm ss}$ represents the ensemble average in the steady state. This phenomenon originates from the metaorder splitters and is well formulated by the Lillo-Mike-Farmer (LMF) model [5, 4]. Particularly, LMF assumed the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$ for the metaorder size $Q$ and predicted the formula
$\gamma=\alpha-1,\>\>\>\alpha\in(1,2),$ (2)
@@
connecting the financial market microstructure and the macroscopic statistics. Recently, the authors validated this prediction quantitatively by scrutinizing the microscopic dataset on the Tokyo Stock Exchange (TSE) [6, 7]. However, this phenomenon is very counter-intuitive: given that the market-order flow is predictable, why is the price dynamics unpredictable? This question is evident given that the price dynamics should exhibit predictable anomalous diffusion if the price impact is linear regarding the metaorder, as documented in the linear propagator model [1, 8, 9] (except under the fine-tuned balance condition between the LRC and the memory). Why is the price dynamics diffusive in the presence of the LRC? This is the first empirical enigma.
@@
The second enigma is the nonlinearity of the price impact, called the square-root law (SRL) [1, 10]. In practice, it has been reported that the price impact is a nonlinear function of the metaorder size $Q$, such that $I(Q):=\langle\epsilon\Delta m\rangle\propto Q^{\delta}$ with $\delta\approx 1/2$, where $\Delta m$ is the price impact by the metaorder $Q$. Recently, the authors scrutinized the TSE dataset, provided a very accurate estimation of $\delta$, and established the strict universality of the SRL, such that $\delta$ exactly equals to $1/2$ for all liquid stocks on the TSE within statistical errors [11]. Still now, the cause of the strict universality of the SRL is unclear. One of the most promising models might be the nonlinear propagator model (called the latent-order book (LLOB) model [1, 10, 12]), but a partially negative evidence was observed on the LLOB hypothesis [13]. Thus, there is no consensus yet regarding the microscopic origin of the SRL.
The third and fourth enigmas are the inverse-cubic law for the price statistics [1, 3, 14, 15, 16, 17, 18] and the volatility clustering [3, 2]: the price movement obeys a fat-tail distribution and the volatility has the long memory:
$P(\Delta m)\propto(\Delta m)^{-\beta-1},\>\>\>\beta\approx 3,$ (3)
$C_{V}(\tau):=$ $\frac{\langle\sigma^{2}(0,t)\sigma^{2}(\tau,t+\tau)\rangle_{\rm ss}}{\langle\sigma^{4}(0,t)\rangle_{\rm ss}}\propto\tau^{-\zeta},\>\>\>\zeta\approx 0.5,$ (4)
where $\sigma^{2}(t,t+\tau):=\{p(t+\tau)-p(t)\}^{2}$. These power laws are ubiquitously observed worldwide independently of asset classes and have been considered as an important part of stylized facts [2, 3]. Despite their long research history, the microscopic origin of these enigmatic laws has been unclear.
@@
In this Letter, we propose a minimal model that naturally extends the LMF framework to incorporate nonlinear price-impact dynamics, thereby unifying these four enigmas from a single, coherent perspective. We consider the system composed of $M$ order splitters whose the metaorder-size statistics obeys the power law $P(Q)\propto Q^{-\alpha-1}$ with $\alpha\in(1,2)$. All traders are assumed to execute their metaorders whose price impact obeys the nonlinear scaling $I(Q)\propto Q^{\delta}$ with $\delta\in(0,1)$. Crucially, this model is exactly solvable. The price-impact contribution by a single trader can be mapped to the Lévy-walk theory [19, 20] with nonlinear walking-speed down [21]. Our exact solution and numerical simulation have three surprising implications: First, the price dynamics is always diffusive even in the presence the LRC if $\delta=1/2$. Second, the price-movement statistics obeys the power law with $\beta=\alpha/\delta$. This result is consistent with the inverse-cubic law because the typical empirical value of $\alpha$ is $3/2$. Third, we numerically find that our model naturally exhibits the volatility clustering. Thus, our simple model is very simple but minimally consistent with the four important enigmas in econophysics with clear exact solutions.
@@
The staring point of our model is the LMF model composed of statistically-independent $M$ order splitters with metaorder size $Q$. The size $Q$ is assumed to obey the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$ with $\alpha\in(1,2)$, and the set of the order splitters is denoted by $\Omega_{\rm TR}:=\{1,2,\dots,M\}$ with a positive integer $M>0$. In this Letter, we keep the essentially-identical setup as for the metaorder splitters.
@@
After finalizing the metaorder execution, trader $i$ randomly resets both order sign $\epsilon^{(i)}=\pm 1$ and metaorder size $Q^{(i)}$, takes a rest according to the resting-time distribution $\psi_{r}(\Delta t)$, and then restart his next metaorder execution. For simplicity, we assume $\nu=1$ and $c=1$ by setting the appropriate time and price units.
In this Letter, we consider this deterministic price-impact case as the minimal assumption, since incorporating Gaussian fluctuations has only a minor qualitative effect. Additionally, the volume $Q$ is assumed to be a real number instead of integers for analytical simplicity.
@@
where $t_{k;\rm ini}^{(i)}$ and $t_{k;\rm fin}^{(i)}$ are the initial and final times of the $k$-th metaorder of trader $i$, and $N^{(i)}(t)$ is the total number of metaorders during $[0,t)$ for trader $i$. Also, the order sign $\epsilon_{k}^{(i)}=\pm 1$ is randomly selected with equal probability, and the final metaorder size $Q_{k}^{(i)}=t_{k;\rm fin}^{(i)}-t_{k;\rm ini}^{(i)}$ obeys the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$. The resting time $\Delta t_{k;\rm rest}:=t_{k+1;\rm ini}^{(i)}-t_{k;\rm fin}^{(i)}$ obeys the exponential resting-time distribution $\psi_{r}(\Delta t)=(1/\tau_{r})e^{-\Delta t/\tau_{r}}$. This is the minimal extension of the LMF model by incorporating the nonlinear price impact, by allowing simultaneous metaorder executions among traders on the continuous time axis.
@@
This model is exactly solvable because the price-impact contribution by a single trader can be mapped to the Lévy-walk theory [19, 20] with nonlinear walking speed [21] (see Appendix A). Indeed, let us define the price-impact contribution by trader $i$ as
@@
Diffusive price dynamics under the square-root law.
Figure 2: (a) Phase diagram between superdiffusion and normal diffusion. The phase boundary is given by $2\delta=\alpha$, implying that the price dynamics is always diffusive by assuming the SRL $\delta=1/2$. (b) Numerical mean-squared displacement for $\delta\in\{0.25,0.5\}$ and $\alpha\in\{1.25,1.5,1.75\}$, showing a consistent behavior with our phase diagram. (c) Numerical mean-squared displacement for $\delta\in\{0.75,1\}$, showing the crossover between superdiffusion (orange markers) and normal diffusion (blue markers). (d) Complementary cumulative distribution function for the price changes, showing the inverse-cubic law $P_{>}(\Delta m):=\int_{\Delta m}^{\infty}P(|\Delta m^{\prime}|){\rm d}\Delta m^{\prime}\propto(\Delta m)^{-\beta}$ with $\beta=3$. We assumed the following parameters: $\delta=1/2$, $\alpha=1.5$, $\tau_{r}=10^{5}$, $t=10^{4}$, and $M=1$. See the inset for the corresponding semi-log plot.
@@
$\langle\Delta m^{2}(t)\rangle\propto\begin{cases}t^{1+2\delta-\alpha}&\mbox{if$2\delta>\alpha$}\\ t&\mbox{if $2\delta<\alpha$}\end{cases},$ (7)
implying that superdiffusion arises if and only if $2\delta>\alpha$. Under the standard LMF assumption $\alpha\in(1,2)$, the price dynamics always exhibits normal diffusion for $\delta\leq 1/2$ (see Fig. 2 (a-c)). See Appendix B for the details.
This result is surprising because it guarantees price diffusion even in the presence of the LRC, assuming the SRL $\delta=1/2$. In other words, thanks to the sufficiently concave nature of the SRL, the market exhibits strong resilience against large liquidity consumption by order splitters. This interpretation is crucially important: it is the SRL that suppresses the large price movement in the presence of the LRC. This scenario highlights the importance of studying the microscopic origin of the SRL for stable regulation of financial markets.
Inverse-cubic law.
We next study the PDF of the price impact $\Delta m$. By evaluating the asymptotic tail of the PDF, we obtain
$P(\Delta m)\propto(\Delta m)^{-\beta-1},\>\>\>\beta:=\frac{\alpha}{\delta}$ (8)
up to the cutoff. This result is clearly consistent with the inverse-cubic law for $\delta=1/2$ because the typical value of $\alpha$ is $3/2$ according to Ref. [5, 22, 6, 7] (see Fig. 2(d)).
For $\beta\leq 2$, the tail can be widely observed for large $t$ because the generalized central limit theorem applies (see Appendix C for its derivation based on the large-deviation principle [23]). On the other hand, for $\beta>2$, the power-law tail regime might not be easily observed depending on the model parameters, because the power-law tail is expected to gradually shrink for large $t$ due to the conventional central limit theorem. However, we identify the parameter regime whereby the power-law tail is clearly observed even for $\beta>2$: When the average resting time satisfies $\tau^{*}\gtrsim t$ with the observation time $t$, we can clearly observe the power-law regime. See Appendix D for the asymptotic evaluation based on the assumption $\tau^{*}\gtrsim t$ (see also Appendix E for a simplified heuristic derivation). This parameter regime is realistic; conversely, it is unrealistic to assume that a splitter rapidly buys and sells huge volumes of stocks within five minutes, given that the inverse cubic law was reported for $t\approx 5$ mins. in [16].
Previously, several theoretical models have been proposed to explain the inverse-cubic law. The first explanation was based on a traditional economic theory [14, 15], one of the main predictions of which was rejected in our previous Letter recently [11]. Some researchers argue that price dynamics follow the Kesten process due to trend-following behavior (see the dealer model [24, 25, 26]), while others suggest that the self-exciting nature of order flows plays a crucial role (see the nonlinear Hawkes process [27, 28, 29], essentially belonging to a nonlinear Kesten family). However, a clear justification has been missing for why $\beta\approx 3$ typically. While our model does not incorporate realistic intraday patterns or external events—both of which we consider essential for accurately predicting $\beta$—it offers a simple and plausible mechanism for why $\beta\approx 3$ is commonly observed.
Volatility clustering.
Figure 3: Volatility ACF in our model, numerically showing the long memory $C_{V}(\tau)\propto\tau^{-\zeta}$ with $\zeta\in(0,1)$. The model parameter was given by $M=1$, $\delta=0.5$, $\tau_{r}=10$, and $t=1000$. Empirically, the power-law exponent obeys $\zeta\approx\alpha-1$.
Furthermore, we numerically find that this model exhibits the volatility clustering (see Fig. 3): i.e., the ACF of the volatility has the long memory, such that
$C_{V}(\tau)\propto\tau^{-\zeta},\>\>\>\zeta\in(0,1),$ (9)
by assuming the SRL $\delta=1/2$ and $\alpha\in(1,2)$. While we have no mathematical derivation yet, we numerically conjecture an empirical relation
$\zeta\approx\alpha-1.$ (10)
We leave its mathematical derivation for future works.
@@
We develop an exactly solvable model of nonlinear price-impact dynamics in the presence of metaorder splitters. By mapping this model to continuous-time random walks, we analytically derive the statistics of the price dynamics. Assuming the square-root law, our exact solutions demonstrate that the price dynamics remains diffusive, even though the order flow is easily predictable. Furthermore, this model naturally exhibits both inverse-cubic law and volatility clustering. These results are surprising, given that our model, which relies solely on the plausible assumptions of metaorder splitting and the SRL, lacks any trivial mechanisms for replicating both inverse-cubic law and volatility clustering. It is a minimal model, free of artificial memory functions or time-dependent external parameters. While our model is very simple, it provides a unified view on the complex financial price dynamics through the exact solution. It would be interesting to scrutinize the exact solutions by extending our model toward more realistic setups, by introducing (i) stochastic price-impact contributions during metaorder splitting and (ii) the heterogeneity of agents.
Here we discuss the implication of our theory. First, our model is available for various numerical statistical estimation regarding the square-root law. Actually, our model was essentially inspired by the numerical statistical model introduced in our previous Letter [11]. In Ref. [11], we estimated the statistical errors of the estimated $\delta$ by studying a numerical price-impact model exactly obeying the SRL: (i) The price dynamics is essentially identical to the rule (5) by adding stochastic contributions. Other conditions are based on the TSE dataset, such that (ii) the metaorder size $Q_{k}^{(i)}$, the starting time $t_{k;\rm ini}^{(i)}$, and the final time $t_{k;\rm fin}^{(i)}$ of the metaorder executions are identical to the TSE dataset, and (ii) the order signs $\epsilon_{k}^{(i)}$ are randomly shuffled to repeat Monte Carlo simulations. We used such a model to numerically study the consistency and unbiasedness of the statistical estimators therein. While we noticed that the the price dynamics in the numerical model was diffusive at that time, its clear reason was elusive. This Letter provides the theoretical reason why our numerical statistical model provided a plausible time-series even in the presence of the metaorder splitters.
@@
Second, our theory clarifies the practical importance of the SRL regarding the market stability. The presence of the LRC implies that markets suffer from the large liquidity consumption by order splitters. It has been mysterious why markets operate consistently with the unpredictable diffusive nature even in the presence of such a large demand. Our theory shows that the concavity of the SRL alleviates the large liquidity consumption by order splitters to lead the diffusive price dynamics. In other words, if the price impact were less concave than the SRL, the price should have been superdiffusive with clear predictability. As a next step, it would be an interesting to study the statistical characters of the coefficient $c$ in the SRL $I(Q)\propto c\sqrt{Q}$ because it characterises the market resilience against the large demand by large investors.
Acknowledgements.
We thank for the fruitful comments by J.-P. Bouchaud and M. Guillaume. YS was supported by JSPS KAKENHI (Grant No. 24KJ1328). KK was supported by JSPS KAKENHI (Grant Nos. 21H01560 and 22H01141) and the JSPS Core-to-Core Program (Grant No. JPJSCCA20200001). YS developed the numerical program code and contributed to the analytical calculations for the mean-squared displacement. KK designed and managed the project, and contributed the analytical calculations, particularly for the inverse-cubic law. Both YS and KK contributed to writing the manuscript and approved its final content.
Appendix A Appendix

### v3

# Why do financial prices exhibit Brownian motion despite predictable order flow?

@@ Abstract
In financial market microstructure, there are two enigmatic empirical laws: (i) the market-order flow has predictable persistence due to metaorder splitters by institutional investors, well formulated as the Lillo-Mike-Farmer model. However, this phenomenon seems paradoxical given the diffusive and unpredictable price dynamics; (ii) the price impact $I(Q)$ of a large metaorder $Q$ follows the square-root law, $I(Q)\propto\sqrt{Q}$. Here we theoretically reveal why price dynamics follows Brownian motion despite predictable order flow by unifying these enigmas. We generalize the Lillo-Mike-Farmer model to nonlinear price-impact dynamics, which is mapped to an exactly solvable Lévy-walk model. Our exact solution shows that the price dynamics remains diffusive under the square-root law, even under persistent order flow. This work illustrates the crucial role of the square-root law in mitigating large price movements by large metaorders, thereby leading to the Brownian price dynamics, consistently with the efficient market hypothesis over long timescales.
econophysics, market microstructure, the square-root law, the Lillo-Mike-Farmer model, Lévy walks
Figure 1: Order-splitting behaviour (left) and the corresponding price impact (right). Institutional traders typically split their large metaorders into a long sequences of small child orders to minimize their transaction cost. The long-term predictability (long memory) of market-order flow arises due to this order splitting because the order signs of the child orders are identical. The resulting market impact—the price change between the initial and final times of metaorders—empirically follows a nonlinear scaling $I(Q)\propto\sqrt{Q}$, called the square-root law.
@@ Introduction.
There are two empirical enigmas in financial market microstructure [1, 2, 3]. The first enigma is the long memory of the market-order flow. A market order is an immediate decision to buy or sell stocks, which is represented by $\epsilon_{t}=+1$ ($\epsilon_{t}=-1$) for the buy (sell) market order at time $t$ (e.g., $\{\epsilon_{t}\}_{t\geq 1}=\{+,+,+,-,+\}$ signifies three buys, one sell, and one buy orders). The long memory implies that such order signs are easily predictable and, thus, the autocorrelation (ACF) of the market-order signs decays very slowly:
[…]
where $E[\dots]$ represents the ensemble average in the steady state and the ACF’s sum diverges as $\sum_{\tau=1}^{\infty}E[\epsilon_{t}\epsilon_{t+\tau}]=\infty$. This formula implies that the order flow has persistence for a very long time, such as for a few days [1].
Then, why does this predictability arise ubiquitously in financial markets? This order-sign predictability originates from the metaorder splitters and is well formulated by the Lillo-Mike-Farmer (LMF) model [4, 5]. The LMF theory argues that institutional investors have hidden large orders (called metaorders $Q$), which are split into a series of child orders. Since the order signs of the child orders should be identical, the order flow has natural predictability (see Fig. 1). Particularly, the LMF theory predicted a formula connecting the market microstructure and the ACF:
[…]
where the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$ is assumed for the metaorder size $Q$. Recently, the authors validated this prediction quantitatively by scrutinizing the microscopic dataset of the Tokyo Stock Exchange (TSE) [6, 7] and, thereby, have established the LMF theory as the causality theory for the long memory.
@@
However, this phenomenon is very counter-intuitive: given that the market-order flow is predictable, why is the price dynamics unpredictable? This puzzle arises because the price dynamics would exhibit predictable anomalous diffusion if the price impact were linear regarding the metaorder size, as documented in the linear propagator model [1, 8, 9] (except under the fine-tuned balance condition between the order-flow and price-impact decays). Why is the price dynamics diffusive despite the predictable order flow? This is the first empirical enigma, particularly because most traditional economic theory predicts linear price-impact models [10].
The second enigma is the nonlinearity of the price impact, called the square-root law (SRL) [1, 11] (see Fig. 1). In practice, the linear price-impact model is valid only for small metaorder size $Q$; for large $Q$, it has been reported that the price impact is a nonlinear function of the metaorder size $Q$:
$I(Q):=E[\epsilon\Delta m\mid Q]\propto Q^{\delta},\>\>\>\delta\approx\frac{1}{2},$ (3)
@@
where $\Delta m$ is the price impact by the metaorder $Q$. Recently, the authors scrutinized the TSE dataset, provided a very accurate estimation of $\delta$, and established the strict universality of the SRL, such that $\delta$ exactly equals to $1/2$ for all liquid stocks on the TSE within statistical errors [12]. Still now, the cause of the strict universality of the SRL is unclear. One of the most promising models might be the nonlinear propagator model (called the latent-order book model [1, 11, 13]), but a partially negative evidence was observed on the latent-order book model [14]. Thus, there is no consensus yet regarding the microscopic origin of the SRL.
@@
Here we propose a minimal model that naturally extends the LMF framework to incorporate nonlinear price-impact dynamics, thereby unifying these two enigmas from a single, coherent perspective. We consider the system composed of $M$ order splitters whose metaorder-size statistics follow the power law $\psi_{m}(Q)\propto Q^{-\alpha-1}$ with $21<\alpha<$ 2. All traders are assumed to execute their metaorders whose price impact obeys the nonlinear scaling $I(Q)\propto Q^{\delta}$ with $0<\delta\leq 1$. Crucially, this model is exactly solvable. The price-impact contribution by a single trader can be mapped to the Lévy-walk theory [15, 16] with nonlinear walking-speed down [17]. Our exact solution shows that the price dynamics is always diffusive even in the presence of the long memory if $\delta=1/2$. This result implies that the SRL plays the crucial role in mitigating the price impact due to the huge liquidity consumption by institutional investors. Additionally, we find that this model exhibits consistent behavior with other empirical laws (such as the inverse-cubic law [1, 3, 18, 19, 20, 21, 22] and volatility clustering [3, 2]). Thus, our simple model is very simple but minimally consistent with various empirical enigmas in finance with clear exact solutions.
@@
The staring point of our model is the LMF model composed of statistically-independent $M$ order splitters with metaorder size $Q$. The size $Q$ is assumed to follow the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$ with $21<\alpha<$ 2, and the set of the order splitters is denoted by $\Omega_{\rm TR}:=\{1,2,\dots,M\}$ with a positive integer $M>0$. In this Letter, we keep the essentially-identical setup as for the metaorder splitters.
@@
Upon completing a metaorder, trader $i$ resets both order sign $\epsilon^{(i)}=\pm 1$ and metaorder size $Q^{(i)}$, takes a rest according to the resting-time distribution $\psi_{r}(\Delta t)$, and restart his next metaorder execution. For simplicity, we set $\nu=c=1$ by appropriate choice of time and price units.
In this work, we consider this deterministic price-impact case as the minimal assumption, since incorporating Gaussian fluctuations has only a minor qualitative effect. Additionally, the volume $Q$ is assumed to be a real number instead of integers for analytical simplicity. Also, the impact decay after metaorder splitting is not considered in this work.
@@
where $t_{k;\rm ini}^{(i)}$ and $t_{k;\rm fin}^{(i)}$ are the initial and final times of the $k$-th metaorder of trader $i$, and $N^{(i)}(t)$ is the total number of metaorders during $[0,t)$ for trader $i$. Also, the order sign $\epsilon_{k}^{(i)}=\pm 1$ is randomly selected with equal probability, and the final metaorder size $Q_{k}^{(i)}=t_{k;\rm fin}^{(i)}-t_{k;\rm ini}^{(i)}$ follows the power-law distribution $\psi_{m}(Q)\propto Q^{-\alpha-1}$. The resting time $\Delta t_{k;\rm rest}:=t_{k+1;\rm ini}^{(i)}-t_{k;\rm fin}^{(i)}$ follows the exponential resting-time distribution $\psi_{r}(\Delta t)=(1/\tau_{r})e^{-\Delta t/\tau_{r}}$. This is the minimal extension of the LMF model by incorporating the nonlinear price impact, by allowing simultaneous metaorder executions among traders on the continuous time axis.
@@
This model is exactly solvable because the price-impact contribution by a single trader can be mapped to the Lévy-walk theory [15, 16] with nonlinear walking speed [17] (see Methods). Indeed, let us define the price-impact contribution by trader $i$ as
@@
Brownian price dynamics under the square-root law.
Figure 3: (a) Phase diagram between superdiffusion and normal diffusion. The phase boundary is given by $2\delta=\alpha$, implying that the price dynamics is always diffusive by assuming the SRL $\delta=1/2$. (b) Numerical mean-squared displacement for $\delta\in\{0.25,0.5\}$, $\alpha\in\{1.25,1.5,1.75\}$, and $N=1$, showing a consistent behavior with our phase diagram. (c) Numerical mean-squared displacement for $\delta\in\{0.75,1\}$, showing the crossover between superdiffusion (orange markers) and normal diffusion (blue markers). More detailed numerical results are presented in Supplementary informationfor other parameters.
@@
$E[\Delta m^{2}(t)]\propto\begin{cases}t^{1+2\delta-\alpha}&\mbox{if $2\delta>\alpha$}\\ t&\mbox{if $2\delta<\alpha$}\end{cases},$ (6)
implying that superdiffusion arises if and only if $2\delta>\alpha$. Under the standard LMF assumption $21<\alpha<$ 2, the price dynamics always exhibits normal diffusion for $\delta\leq 1/2$ (see Fig. 3). See Methods for the details. Note that this result is consistent with the linear propagator model [8, 9] for the specific case with non-zero permanent impact for $\delta=1$.
This result is surprising because it guarantees price diffusion even in the presence of the long memory in the order flow, assuming the SRL $\delta=1/2$. In other words, thanks to the sufficiently concave nature of the SRL, the market exhibits strong resilience against large liquidity consumption by order splitters. This interpretation is crucially important: it is the SRL that suppresses the large price movement in the presence of the long memory. This scenario highlights the importance of studying the microscopic origin of the SRL for stable regulation of financial markets.
@@
We develop an exactly solvable model of nonlinear price-impact dynamics in the presence of metaorder splitters. By mapping this model to continuous-time random walks, we analytically derive the statistics of the price dynamics. Assuming the SRL, our exact solution demonstrates that the price dynamics remains diffusive, even though the order flow is easily predictable. It would be interesting to scrutinize the exact solutions by extending our model toward more realistic setups, by introducing (i) stochastic price-impact contributions during metaorder splitting and (ii) the heterogeneity of agents.
@@
Here we discuss the implication of our theory. First, our theory clarifies the practical importance of the SRL regarding the market stability. The presence of the long memory implies that markets suffer from the large liquidity consumption by order splitters. It has been mysterious why markets operate consistently with the unpredictable diffusive nature even in the presence of such a large demand. Our theory shows that the concavity of the SRL alleviates the large liquidity consumption by order splitters, which leads to diffusive price dynamics. In other words, if the price impact were less concave than the SRL, the price should have been superdiffusive with clear predictability. As a next step, it would be an interesting topic to study the statistical characters of the coefficient $c$ in the SRL $I(Q)\propto c\sqrt{Q}$ because it characterises the market resilience against the large demand by large investors.
Second, our theory highlights that the SRL is not a technical minor topic but should be regarded as a key topic in financial economics because it is directly related to the theoretical foundation of the efficient market hypothesis (EMH)—the hypothesis stating that the market is sufficiently efficient and the price is unpredictable (thus diffusive) from public information. Evidently, the EMH is the most important concept for various financial studies. However, most of traditional economic theories (such as Kyle’s one [10]) predict the linear price-impact law, which apparently contradicts the EMH in the presence of the predictable market-order flow. Our theory clarifies that, even under the predictable market-order flow, the EMH robustly holds at a long timescale thanks to the SRL. In other words, it is the SRL that prohibits simple arbitrage strategies from professional traders even when the market-order flow has the predictable long memory. There is the room to update financial-economic theories by incorporating the SRL to keep the theoretical consistency between the EMH and the predictable market-order flow.
Third, our model is available for various numerical statistical estimation regarding the SRL. Actually, our model was essentially inspired by the numerical statistical model introduced in our previous Letter [12]. In Ref. [12], we estimated the statistical errors of the estimated $\delta$ by studying a numerical price-impact model exactly obeying the SRL: (i) The price dynamics is essentially identical to the rule (4) by adding stochastic contributions. Other conditions are based on the TSE dataset, such that (ii) the metaorder size $Q_{k}^{(i)}$, the starting time $t_{k;\rm ini}^{(i)}$, and the final time $t_{k;\rm fin}^{(i)}$ of the metaorder executions are identical to the TSE dataset, and (ii) the order signs $\epsilon_{k}^{(i)}$ are randomly shuffled to repeat Monte Carlo simulations. We used such a model to numerically study the consistency and unbiasedness of the statistical estimators therein. While we noticed that the the price dynamics in the numerical model was diffusive at that time, its clear reason was elusive. This Letter provides the theoretical reason why our numerical statistical model provided a plausible time-series even in the presence of the metaorder splitters.
Fourth, interestingly, our simple model is consistent with even other enigmatic empirical laws in finance without fine-tuning. For example, the price change $\Delta m$ follows the inverse-cubic law [1, 3, 18, 19, 20, 21, 22]
$P(\Delta m)\propto(\Delta m)^{-\beta-1},\>\>\>\beta:=\frac{\alpha}{\delta},$ (7a)
clearly consistent with the inverse-cubic law $\beta\approx 3$ for $\delta=1/2$ because the typical value of $\alpha$ is $3/2$ according to Ref. [4, 23, 6, 7] (see Methods). Furthermore, our model exhibits even volatility clustering [3, 2]:
$C_{V}(\tau)$ $:=\frac{E[(\sigma^{2}(0,t)-\mu_{\sigma^{2}_{t}})(\sigma^{2}(\tau,\tau+t)-\mu_{\sigma^{2}_{t}})]}{E[(\sigma^{2}(0,t)-\mu_{\sigma^{2}_{t}})^{2}]}$ (7b)
[…]
where $\mu_{\sigma^{2}_{t}}=E[\sigma^{2}(0,t)]$, and $\sigma^{2}(t,t+\tau):=\{p(t+\tau)-p(t)\}^{2}$, by assuming the SRL $\delta=1/2$. While we have no mathematical derivation, we numerically conjecture $\zeta\approx\alpha-1$ (see Methods and Fig. 4). These two characters (7) are surprising, given that our model, which relies solely on the plausible assumptions of metaorder splitting and the SRL, lacks any trivial mechanisms for replicating both inverse-cubic law and volatility clustering. It is a minimal model, free of artificial memory functions or time-dependent external parameters.
Finally, let us address the potential limitation of our theory. Our model incorporates neither market-microstructure dynamics nor time-dependent parameters in our model—be they endogenous (e.g., self-excitation mechanism [24, 25, 26] and the limit-order book dynamics) or exogenous (e.g., intraday seasonality and external shocks)—, which is both a strength and a weakness. A strength is the model’s parameter-free nature, allowing direct testability with microscopic datasets. A weakness is its inability to flexibly capture realistic endogenous/exogenous market microscopic dynamics, and we are unsure yet if our simple model—accounting for only order-splitting and the SRL—fully captures the causal mechanisms behind the inverse-cubic law and the volatility clustering. Several researchers claim that limit-order-book conditions and liquidity crises are key to understanding the large price jump [27, 28], but these points are missing in our model. Incorporating such mechanisms into our model would be an important next step.
I Methods